40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 4 04:51:14 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdfCARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdfGlitch_in_CARM_and_PRCL_3Feb2015.pngslowCM_04-02-2015_042805.png
    Reply  Thu Feb 5 04:35:14 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path 4Feb2015_Transition_CARM_REFL11_CM_SLOW_AP_18.png
       Reply  Fri Feb 6 03:21:17 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path 
Message ID: 10979     Entry time: Thu Feb 5 04:35:14 2015     In reply to: 10971     Reply to this: 10982
Author: diego 
Type: Update 
Category: LSC 
Subject: CARM Transition to REFL11 using CM_SLOW Path 

[Diego, Eric]

Tonight was a sad night... We continued to pursue our strategy, but with very poor results:

  • before doing anything, we made sure we had a good initial configuration: we renormalized the arm powers, retuned the X arm green beatnote, did extensive ASS alignment;
  • since the beginning of the night we faced a very uncooperative PRMI, which caused a huge number of locklosses, often just by itself, without even managing to reduce the MICH offset before reducing the CARM one;
  • we had to reduce the PRCL gain to -0.002 in order to acquire PRMI lock, but keeping it such or restoring it to -0.004 once lock was acquired either didn't improve the PRMI stability at all;
  • we also tweaked a bit the PRCL and MICH UGF servos (namely, their frequencies to ~80 Hz and ~40 Hz respectively) and that seemed to help earlier during the night, but not much longer;
  • we only managed to transition CARM to REFL11 via CM SLOW twice;
    • the first time we lost lock almost immediately, probably because of a non-optimal offset between CARM A and B;
    • the second time we managed to stay there a little longer, but then some spike in the PRCL loop and/or the MICH loop hitting the rails threw us out of lock (see the lockloss plot);
    • both times we transitioned at arm power ~18;
  • during the night we used an increased analog ASDC whitening gain, as from Eric's elog here http://nodus.ligo.caltech.edu:8080/40m/10972 ; even with this fix, though, MICH is still often hitting the rails and causing the lock losses;
  • the conclusion for tonight is that we need to figure what is going on with the PRMI...

 

Attachment 1: 4Feb2015_Transition_CARM_REFL11_CM_SLOW_AP_18.png  429 kB  | Hide | Hide all
4Feb2015_Transition_CARM_REFL11_CM_SLOW_AP_18.png
ELOG V3.1.3-